In ΔABC, the measure of ∠A is 36°. The measure of ∠B is three times the measure of ∠C. Find m∠B & m∠C.
Please help me. I only need this one last problem before I finish my homework! Thanks!

Answer by ankor@dixie-net.com(22740)

In ΔABC, the measure of ∠A is 36°.
The measure of ∠B is three times the measure of ∠C.
Find m∠B & m∠C.
:
A = 36
B = 3C
:
A + B + C = 180
36 + B + C = 180
B + C = 180 - 36
B + C = 144
Replace B with 3C
3C + C = 144
4C = 144
C = 144/4
C = 36 degrees
:
B = 3(36)
B = 108 degrees